# OPeRA: A Dataset of Observation, Persona, Rationale, and Action for Evaluating LLMs on Human Online Shopping Behavior Simulation

### v2

@@ Abstract
Can large language models (LLMs) accurately simulate the next web action of a specific user? While LLMs have shown promising capabilities in generating “believable” human behaviors, evaluating their ability to mimic real user behaviors remains an open challenge, largely due to the lack of high-quality, publicly available datasets that capture both the observable actions and the internal reasoning of an actual human user. To address this gap, we introduce OPeRA, a novel dataset of Observation, Persona, Rationale, and Action collected from real human participants during online shopping sessions [^1] The dataset is currently available upon request.. OPeRA is the first public dataset that comprehensively captures: user personas, browser observations, fine-grained web actions, and self-reported just-in-time rationales. We developed both an online questionnaire and a custom browser plugin to gather this dataset with high fidelity. Using OPeRA, we establish the first benchmark to evaluate how well current LLMs can predict a specific user’s next action and rationale with a given persona and <observation, action, rationale> history. This dataset lays the groundwork for future research into LLM agents that aim to act as personalized digital twins for human.
@@ 1 Introduction
Figure 1: The overall OPeRA data collection process and user action prediction benchmark experiment. We first developed ShoppingFlow plugin (details in Figure 2) and collected 740 user shopping session from 54 unique users in four weeks, resulting in 36,691 real-user <actions, observation> pairs and 721 user-annotated rationales (details in Figure 3). We then benchmarked four LLMs’ performance on the user action prediction tasks, where input includes user persona, and past <actions, rationales, observations> and the output is the next <action, rationale> (results in Table 4 and 5).
@@
Large language model (LLM) agents have exhibited impressive performance across diverse tasks, including planning, reasoning, and acting in web-based environments Xie et al. (2024); Yao et al. (2023); Jin et al. (2025). A promising frontier in this area is human behavior simulation, where LLM agents generate user-like action sequences on digital platforms Chen et al. (2025). These agents (i.e., role-playing agents) are increasingly used in applications such as UI/UX testing Lu et al. (2025b), social science research Park et al. (2024), accessibility testing Taeb et al. (2024), and personal digital assistants OpenAI (2025b). Yet, while these agents can generate believable behavior, the more critical question remains: can they generate behavior that is accurately aligned with real human?
@@
Despite progress in agent-based simulation Park et al. (2023, 2024), most existing evaluations focus on aggregate outcomes (e.g., survey responses or end-task completions) while overlooking the step-wise rationale and actions that underlie user behavior pattern Chen et al. (2025). For example, Park et al. (2024) compared LLM agents’ survey results with real humans by replicating various social science studies. Furthermore, many role-play agents rely solely on prompting without grounding in real human data training, which limits their accuracy and personalization. Although some recent efforts incorporate user behavior data via fine-tuning Lu et al. (2025a), these datasets are often proprietary or lack critical detail, such as the reasoning or persona behind user actions.
@@
Current open-source datasets for user behavior simulation fall short in several key aspects. First, most datasets record only sparse, decontextualized–or even synthetic–user actions. Some shopping datasets like Amazon-M2 or ECInstruct Jin et al. (2024a, b) record only the isolated actions (e.g., purchases or clicks) with limited observation context, while others Deng et al. (2024); Chen et al. (2024); Yao et al. (2022) use synthetic or third-party annotated behaviorial data, which lacks the individual behavior pattern and the authenticity. Additionally, few datasets provide step-wise reasoning or persona information, despite prior work has shown that rationale can improve LLM agent’s performance in behavior and decision modeling Lu et al. (2025a); DeepSeek-AI et al. (2025); and user persona strongly correlates with behavioral patterns Helmi et al. (2023), making persona essential for durable personalization experience.
@@
To address these limitations, we introduce OPeRA–a dataset of Observation, Persona, Rationale, and Action collected from real human users during online shopping. OPeRA provides rich, time-aligned logs of users’ web browsing behavior, completed with self-reported rationales and detailed self-reported persona profiles. Unlike prior datasets, OPeRA captures not only what users do but also why they do it, enabling deeper insights into decision-making processes. This paper focus on the online shopping domain as a beginning point due to its everyday prevalence, complex decision flow, and strong ties to personalization Mican and Sitar-Taut (2020); Wei (2016). Online e-commerce environments like Amazon requires a user’s multi-step interactions involving comparisons, trade-offs, and goal-directed behaviors in one shopping session, all of which are prime testbeds for studying the capacity of LLMs to simulate real human actions.
@@
With 740 shopping sessions from 54 unique users, 36,691 <observation, action> pairs, and 721 human-annotated rationales, OPeRA serves as the first benchmark dataset for evaluating LLM agents on personalized and verifiable user behavior simulation. We benchmark four state-of-the-art LLMs (GPT-4.1 OpenAI (2025a), DeepSeek-R1 DeepSeek-AI et al. (2025), Claude-3.7 Anthropic (2025), and Llama-3.3 Meta (2024)) and analyze their ability to predict the next action and rationale of a specific user based on their persona and interaction history. These findings lay a foundation for future work on building LLM-powered digital twins capable of accurate and adaptive behavior modeling.
@@
Table 1: Properties of existing datasets compared to OPeRA. “O”: Environment Observation. “Pe”: User Persona. “R”: Rationale behind action. “A”: Action Space. “Source”: Action Source. OPeRA contains 740 sessions, which includes rationales, user personas, and 36691 <observation, action> pairs from real users.
@@ 2.1 LLMs for Human Behavior Simulation
Large Language Model agents can handle complex tasks Wang et al. (2024c); Yao et al. (2023); Shinn et al. (2023); Wu et al. (2023), and researchers are using them as human proxies across domains, from social-science simulations Park et al. (2023); Wang et al. (2024a); Sreedhar et al. (2025) and recommender-system evaluation Wang et al. (2023) to UX testing Lu et al. (2025b) and health-counsellor training Louie et al. (2025). They even reproduce classic results in experimental psychology and economics, such as Milgram Shock Experiment Aher et al. (2023).
@@ 3 OPeRA Dataset
Our work presents a novel dataset OPeRA to advance NLP research in realistic user behavior simulation, particularly within the context of online shopping scenarios. Unlike existing resources that lack step-wise user behavior trajectories or rationales behind user behaviors, OPeRA comprises of $740$ authentic shopping sessions on Amazon.com conducted by $54$ unique customers.
@@ 3.2 Dataset Details
An overview of the dataset is presented in Figure 3. Of all the users, 54 contributed at least one shopping session. In total, the dataset contains 740 sessions, 36,691 <observation, action> pairs, and 721 rationale annotations. Table 2 presents the core statistics of user actions and rationales in OPeRA.
@@
Click 9,339 (25.45%) 297 (41.19%) 5.90
Tab Activate 2,226 (6.07%) 63 (8.74%) 5.89
Navigate 2,223 (6.06%) 32 (4.44%) 5.56
Text Input 814 (2.22%) 0 –
@@
Table 2: Full dataset statistics. “Len” reports the average rationale length (in words) for each action type.
4 Tasks and Experiments
4.1 Tasks
We show how OPeRA can be leveraged to evaluate LLM’s ability to simulate consumer behavior in online shopping. The following two tasks are used to evaluate models’ performance on user behavior simulation with inputs including user persona $P_{i}$, action traces $A_{j}$, web context $O_{j}$ and rationale $R_{j}$, for participant $i$ and session $j$.
4.1.1 Next Action Prediction
$a_{t}=F_{action}(a_{1...t-1},r_{1...t-1},o_{1...t},P_{i})$
4.1.2 Joint Rationale and Action Generation
The reasoning process plays an important role in reflecting how well models understand user behavior Wang et al. (2024b). This task aims to jointly generate a textual rationale $r_{t}$, which describes why the user would perform an action at step $t$, and the corresponding action $a_{t}$.
$<a_{t},r_{t}>=F_{\text{joint}}(a_{1...t-1},r_{1...t-1},o_{1...t},P_{i})$
4.2 Experiments
4.2.1 Data Processing
To adapt the raw user behavior data to the behavior simulation task, we applied several processing steps prior to input.
Action Space Simplification
To support behavior modeling and evaluation under a more tractable setting, we follow prior work Lu et al. (2025a) and similarly define a simplified action $\mathcal{A}$ space consisting of key interaction types. Specifically, we retain four high-level actions that are both semantically meaningful and commonly observed: input, click, purchase, and terminate. Within the “click” category, we further differentiate between three subtypes to reflect distinct user intents: clicking on a filter option, clicking the search box, and clicking on other page elements, primarily product items. This abstraction reduces complexity while preserving the core structure of user behaviors, enabling more stable behavior simulation.
Filtering
To reduce noise and improve the quality of behavioral data, some actions that do not reflect meaningful intent are removed. Specifically, we filtered out clicks on non-interactive areas such as the background, and further filtered actions involving Amazon Rufus, resulting in 715 sessions after filtering. The session action distribution after processing are shown in Table  3. Appendix G further shows total action distribution.
@@ 4.2.1 Data Processing
Avg. # of Action 13.64
Avg. # of Input 0.92
Avg. # of Click 11.72
Avg. # of Purchase 0.45
Avg. # of Termination 0.55
Avg. # of Rationale 0.39
@@
Table 3: Summary of processed dataset statistics
@@ 4.2.2 Experiment Setup
The test set is constructed by randomly sampling 15 out of 54 users and randomly sampling 100 sessions from these users, resulting in 1433 actions and 50 rationales. We evaluate four state-of-the-art LLMs, including two open-source models, Llama-3.3-70B-Instruct and DeepSeek-R1, and two proprietary models, GPT-4.1 and Claude-3.7-Sonnet.
All models are evaluated in a zero-shot, prompt-based setting without fine-tuning. Prompt templates are provided in Appendix H.
@@
4.2.3 Evaluation
To assess the accuracy of predicted user actions, we apply an exact match criterion: a prediction is correct only when all required components align with the ground truth. Specifically, for click actions, both the specific subtype (e.g., clicking on filter, search box, and others) and the clicked target must match. For input actions, this includes identifying the input field and generating exact input text.
The quality of generated rationales is evaluated by comparing them with the ground-truth rationales using the BERTScore F1 Zhang et al. (2019), and the text similarity metrics ROUGE-L Lin (2004).
4.2.4 Results and Analysis
The results for next action prediction are presented in Table 4. Evaluation results for the joint rationale and action generation task are provided in Table 5.
@@ 4.2.4 Results and Analysis
Table 4: Evaluation of actions in next action prediction task. “Claude-3.7”: Claude-3.7-Sonnet, “Llama-3.3”: Llama-3.3-70B-Instruct. All metrics are reported as percentages (%). Data sample size $n=1,333$.
{booktabs}
[…]
   w/o persona 27.08 69.10 84.31 10.05
      w/o rationale 22.92 63.90 84.39 10.56
[…]
   w/o persona 16.67 77.73 83.00 9.48
      w/o rationale 16.67 66.02 83.40 9.14
[…]
   w/o persona 20.83 58.87 84.72 7.94
      w/o rationale 16.67 39.32 84.57 9.41
[…]
   w/o persona 14.58 56.67 61.31 9.68
      w/o rationale 10.42 61.45 66.74 9.20
[…]
Table 5: Joint evaluation of rationale and action prediction. “Llama-3.3-70B”: Llama-3.3-70B-Instruct. All metrics are reported as percentages (%). Instance size $n=48$. For generated action, we evaluate the accuracy and the action type F1. For generated rationale, we evaluate with BERTScore and ROUGE-L.
Next Action Prediction
Table 4 reports exact-match accuracy for action generation and F1 for action type classification.
Specifically, GPT-4.1 substantially outperforms other models, achieving an 18.53% exact match accuracy and a 68.40% F1 score. This strong performance may be attributed to its larger context window, which facilitates better processing of complex interaction histories. Claude-3.7-Sonnet exhibits moderate performance, while both DeepSeek-R1 and Llama-3.3-70B-Instruct underperform. This is possibly due to their limited 128k context length, which may hinder their ability to capture long-range dependencies in interaction histories.
The role of persona information varies across models. GPT-4.1 effectively leverages persona data, showing clear performance degradation without it. Conversely, other models demonstrate comparable or even improved performance without persona information, suggesting they struggle to effectively model personalized user behavior.
Additionally, removing the history rationale generally reduces model performance, highlighting its importance for predicting next actions.
Joint Rationale and Action Generation
As shown in Table 5, the generated rationale demonstrates relatively low textual similarity metrics compared to ground-truth. Besides, the inconsistent performance patterns across models indicate instability in rationale generation capabilities.
Table 5 also indicates a performance improvement in action prediction when models generate both rationale and actions simultaneously. GPT-4.1 leads in exact match accuracy (27.08%), while Claude-3.7-Sonnet achieves the highest action type F1 score (83.25%). Deepseek-R1 performs moderately, whereas Llama-3.3-70B-Intruct continues to underperform, further emphasizing its limitations with handling complex rationale and action prediction sequences. Notably, overall performance is higher in the joint generation setting compared to the standalone action prediction task, suggesting that explicitly generating rationales alongside actions enhances performance.
For persona information, GPT-4.1 and Claude-3.7-Sonnet leverage user persona well, showing increasing performance. On the other hand, DeepSeek-R1 and Llama-3.3-70B-Instruct consistently struggle in dealing with this information.
The importance of rationale is further supported by the consistent performance degradation observed across all models when history rationales are removed.

### v4

@@ Abstract
Can large language models (LLMs) accurately simulate the next web action of a specific user? While LLMs have shown promising capabilities in generating “believable” human behaviors, evaluating their ability to mimic real user behaviors remains an open challenge, largely due to the lack of high-quality, publicly available datasets that capture both the observable actions and the internal reasoning of an actual human user. To address this gap, we introduce OPeRA, a novel dataset of Observation, Persona, Rationale, and Action collected from real human participants during online shopping sessions. OPeRA is the first public dataset that comprehensively captures: user personas, browser observations, fine-grained web actions, and self-reported just-in-time rationales. We developed both an online questionnaire and a custom browser plugin to gather this dataset with high fidelity. Using OPeRA, we establish the first benchmark to evaluate how well current LLMs can predict a specific user’s next action and rationale with a given persona and <observation, action, rationale> history. This dataset lays the groundwork for future research into LLM agents that aim to act as personalized digital twins for human  [^1] https://huggingface.co/datasets/NEU-HAI/OPeRA.
@@ 1 Introduction
Figure 1: We developed ShoppingFlow plugin (Figure 2) to collect user shopping behavior over a four-week period, resulting in OPeRA-full dataset. This dataset comprises 692 sessions from 51 unique users, containing 28,904 real-user <action, observation> pairs and 604 user-annotated rationales (Figure 3). After postprocessing, we obtained OPeRA-filtered, which includes 527 sessions, 5,856 <action, observation> pairs, and 207 rationales. We then benchmarked four LLMs’ performance on user next action prediction task, results in Table 6).
@@
Large language model (LLM) agents have exhibited impressive performance across diverse tasks, including planning, reasoning, and acting in web-based environments Xie et al. (2024); Yao et al. (2023); Jin et al. (2025). A promising frontier in this area is human behavior simulation, where LLM agents generate user-like action sequences on digital platforms Chen et al. (2025). These agents (i.e., role-playing agents) are increasingly used in applications such as UI/UX testing Lu et al. (2025b), social science research Park et al. (2024), accessibility testing Taeb et al. (2024), and personal digital assistants OpenAI (2025b). Yet, while these agents can generate believable behavior, the more critical question remains: can they generate behavior that accurately aligns with real human?
@@
Despite progress in agent-based behavior simulation Park et al. (2023, 2024), current works still have limitations. First, most existing evaluations focus on aggregate outcomes (e.g., survey responses or end-task completions). These methods overlook the step-wise rationale and actions that underlie user behavior patterns Chen et al. (2025). For example, Park et al. (2024) compared LLM agents’ survey results with real humans by replicating various social science studies. Furthermore, many role-play agents rely solely on prompting without grounding in real human data training, which limits their accuracy and personalization. Although some recent efforts incorporate user behavior data via fine-tuning Lu et al. (2025a), these datasets are often proprietary or lack critical detail, such as the reasoning or persona behind user actions.
@@
Current open-source datasets for user behavior simulation fall short in several key aspects. First, most datasets record only sparse, decontextualized–or even synthetic–user actions. Some shopping datasets like Amazon-M2 or ECInstruct Jin et al. (2024a, b) record only the isolated actions (e.g., purchases or clicks) with limited observation context. Others Deng et al. (2024); Chen et al. (2024); Yao et al. (2022) use synthetic or third-party annotated behaviorial data, which lacks the individual behavior pattern and the authenticity. Additionally, few datasets provide step-level reasoning or persona information, despite prior work has shown that rationale can improve LLM agent’s performance in behavior and decision modeling Lu et al. (2025a); DeepSeek-AI et al. (2025). Similarly, user persona strongly correlates with behavioral patterns Helmi et al. (2023), making persona essential for durable personalization experience.
@@
To address these limitations, we introduce OPeRA, a dataset of Observation, Persona, Rationale, and Action collected from real human users during online shopping. OPeRA provides rich, time-aligned logs of users’ web browsing behavior, completed with self-reported rationales and detailed self-reported persona profiles. Unlike prior datasets, OPeRA captures not only what users do but also why they do it, enabling deeper insights into decision-making processes. This paper focus on the online shopping domain as a beginning point due to its everyday prevalence, complex decision flow, and strong ties to personalization Mican and Sitar-Taut (2020); Wei (2016). Online e-commerce environments like Amazon requires a user’s multi-step interactions involving comparisons, trade-offs, and goal-directed behaviors in one shopping session, all of which are prime testbeds for studying the capacity of LLMs to simulate real human actions.
@@
The OPeRA-full contains 692 shopping sessions from 51 unique users, 28,904 <action, observation> pairs, and 604 human-annotated rationales. After post-processing, we also provide OPeRA-filtered, which includes 527 sessions, 5856 <observation, action> pairs and 207 rationales. OPeRA serves as the first benchmark dataset for evaluating LLM agents on personalized and verifiable user behavior simulation. We benchmark four state-of-the-art LLMs (GPT-4.1 OpenAI (2025a), DeepSeek-R1 DeepSeek-AI et al. (2025), Claude-3.7 Anthropic (2025), and Llama-3.3 Meta (2024)) on OPeRA-test (a subset of the OPeRA-filtered) and analyze their ability to predict the next action and rationale of a specific user based on their persona and interaction history. These findings lay a foundation for future work on building LLM-powered digital twins capable of accurate and adaptive behavior modeling.
@@
Table 1: Properties of existing datasets compared to OPeRA. “O”: Environment Observation. “Pe”: User Persona. “R”: Rationale behind action. “A”: Action Space. “Source”: Action Source.
@@ 2.1 LLMs for Human Behavior Simulation
Large Language Model agents can handle complex tasks Wang et al. (2024b); Yao et al. (2023); Shinn et al. (2023); Wu et al. (2023), and researchers are using them as human proxies across domains, from social-science simulations Park et al. (2023); Wang et al. (2024a); Sreedhar et al. (2025) and recommender-system evaluation Wang et al. (2023) to UX testing Lu et al. (2025b) and health-counsellor training Louie et al. (2025). They even reproduce classic results in experimental psychology and economics, such as Milgram Shock Experiment Aher et al. (2023).
@@ 3 OPeRA Dataset
Our work presents a novel dataset OPeRA to advance NLP research in realistic user behavior simulation, particularly within the context of online shopping scenarios.
@@ 3.1.4 Post-Processing
To reduce noise and improve the quality of behavioral data, some actions that occurred on uncommon or rarely visited pages or do not reflect meaningful intent are removed. Moreover, we filtered out clicks on non-interactive areas such as the background, and further filtered actions involving Amazon Rufus.
To support behavior modeling and evaluation under a more tractable setting, we follow prior work Lu et al. (2025a) and similarly define a simplified action $\mathcal{A}$ space consisting of key interaction types. Specifically, we retain three high-level actions that are both semantically meaningful and commonly observed: input, click, and terminate. Within the “click” category, we further differentiate between several subtypes to reflect distinct user intents, shown in Table  3.2. This abstraction reduces complexity while preserving the core structure of user behaviors, enabling more stable behavior simulation.
@@ 3.2 Dataset Details
An overview of the dataset is presented in Figure 3. Of all the users, 51 contributed at least one shopping session. In total, the OPeRA-full contains 692 sessions, 28,904 <action, observation> pairs, and 604 rationale annotations. The OPeRA-filtered contains 527 sessions, 5,856 <action, observation> pairs, and 207 rationales. Table 2 presents the statistics. Table 3.2 shows the click type distribution.
@@
Click 5,253 (18.1%) 5,051 (86.3%)
Tab Activate 1,945 (6.7%) –
Navigate 1,901 (6.6%) –
Text Input 606 (2.1%) 597 (10.2%)
Terminate – 208(3.6%)
@@
Table 2: Action type distribution.
{booktabs}
[…]
review 1052 20.8%
search 763 15.1%
product_option 700 13.9%
product_link 537 10.6%
other 449 8.9%
purchase 321 6.4%
nav_bar 283 5.6%
page_related 198 3.9%
quantity 191 3.8%
suggested_term 182 3.6%
cart_side_bar 145 2.9%
cart_page_select 139 2.8%
filter 91 1.8%
[…]
Table 3: Click type distribution in OPeRA-filtered. Detailed descriptions in Appendix G
@@
Avg. # of Action 11.11
Avg. # of Input 1.13
Avg. # of Click 9.58
Avg. # of Terminate 0.39
@@
Table 4: OPeRA-filtered dataset statistics per-session
4 Tasks and Experiments
4.1 Tasks
We show how OPeRA can be leveraged to evaluate LLM’s ability to simulate consumer behavior in online shopping.
Next Action Prediction
$a_{t}=F_{action}(a_{1...t-1},r_{1...t-1},o_{1...t},P_{i})$
4.2 Experiments
@@ 4.2.1 Experiment Setup
From the OPeRA-filtered, we furthur construct the test set OPeRA-test by randomly sampling 15 out of 51 users and randomly sampling 90 sessions from these users, resulting in 992 actions. We evaluate four state-of-the-art LLMs, including two open-source models, Llama-3.3-70B-Instruct and DeepSeek-R1, and two proprietary models, GPT-4.1 and Claude-3.7-Sonnet.
All models are evaluated in a zero-shot, prompt-based setting without fine-tuning. Prompt templates are provided in Appendix I.
@@
Table 5: Evaluation of actions in next action prediction task. We report four metrics here to assess model performance (Full results can be found in Appendix H): Action Generation Accuracy measures the exact-match accuracy of the predicted next action; Action Type Macro F1 evaluates the model’s ability to predict the correct high-level action category (e.g., input, click, terminate); Click Type Weighted F1 captures the performance of predicting the specific type of click actions; Session Outcome Weighted F1 reflects how well the model can predict the final outcome of a session, where each session ends either in a purchase or a terminate action. “Claude-3.7”: Claude-3.7-Sonnet, “Llama-3.3”: Llama-3.3-70B-Instruct. All metrics are reported as percentages (%). Instance size $n=902$.
4.2.2 Evaluation
To assess the accuracy of generated user actions, we apply an exact match criterion: a prediction is correct only when all required components align with the ground truth. Specifically, for click actions, the clicked target name must match. For input actions, this includes identifying the input field and generating exact input text.
In addition to exact match accuracy, we evaluate the model’s ability to classify action types. We report weighted F1 for high-level action types (click, input, terminate). Given the highly imbalanced nature of user behavior distributions, we also report macro F1 to highlight performance across all classes regardless of frequency.
To further examine fine-grained prediction capabilities, we evaluate the weighted F1 score for click subtypes. This captures whether the model not only predicts that a user will click but also understands the specific type of click (e.g., review, product_link, purchase).
Finally, given the goal-oriented nature of online shopping, we assess the model’s ability to predict session outcomes. Each session in the ground truth ends in either a click on purchase-related button or a terminate action. We evaluate the model’s accuracy and F1 score on these terminal actions to understand whether it can correctly capture users’ long-term goals and decision-making processes.
4.2.3 Results and Analysis
The results for next action prediction are presented in Table 6. Among all models, GPT-4.1 achieves the strongest overall performance. It obtains the highest action generation accuracy of 22.06%, a macro F1 score of 48.78% on action type prediction, and a weighted F1 score of 44.47% on click type prediction. For session outcome prediction, GPT-4.1 also shows solid performance with 58.47% F1. This strong performance may be attributed to its large context window, which facilitates better processing of complex interaction histories. DeepSeek-R1 performs moderately well across tasks. It achieves a maximum action generation accuracy of 15.74% and obtains solid outcome prediction, possibly due to its strong reasoning abilities. However, the performance may be limited by the model’s 128k context length. Claude-3.7 shows modest performance. Its action generation accuracy is around 10.75%, and action type prediction is generally weaker than GPT-4.1 and DeepSeek. Nonetheless, it achieves relatively good outcome prediction, suggesting some robustness in capturing high-level user intent. LLaMA-3.3 underperforms across all metrics, with action generation accuracy of only 8.76%. Its lower performance may be due to the smaller model size (70B) and shorter context length (128k).
The role of persona information varies across models. While adding persona information does not consistently improve exact action generation accuracy, it generally enhances the model’s performance on action type and click type prediction across all model families. This suggests that persona information provides useful priors about user preferences and behavior patterns, helping the model better classify action semantics. However, its limited effect on action generation accuracy implies that simply including persona in the prompt may introduce noise rather than help. Current models have limited ability to deeply integrate persona into step-level decision-making. This highlights potential room for improvement in personalized user modeling.
Additionally, removing historical rationales consistently leads to performance degradation across most models and metrics, particularly in outcome prediction. This confirms that rationale information serves as valuable intermediate supervision, guiding the model to align its decisions with plausible user intent. We also note several outliers in the results. For instance, LLaMA-3.3 does not consistently benefit from rationale inputs. This may be due to its smaller model size and limited capacity to leverage additional contextual signals effectively.